# ASVD: Activation-aware Singular Value Decomposition for Compressing Large Language Models

### v1

@@ 4.4 Decomposed Network Analysis
In this section, we conduct a detailed analysis of the decomposed network. Figure 5 presents the per-type parameters ratio and per-block parameters ratio. Observing the plot, we note that parameters in the MLP components (gate projection, up projection, and down projection) exhibit minimal compression. In MHA, the V projection layer experiences relatively small compression, whereas q projection and k projection can be significantly compressed, indicating redundancy in these components. Turning our attention to the per-block compression ratio, we find that the first layer can undergo substantial compression. In contrast, the compression ratios for the other layers, except for two middle layers, show similar compression rates.
@@ 5 Conclusion
This study introduces a training-free approach to compress the Large Language Models (LLMs), enhancing their usability across diverse computing environments. The Activation-aware Singular Value Decomposition (ASVD) and Sensitivity-based Truncation Rank Searching (STRS), central to our approach, adeptly addresses activation outliers and varying layer sensitivities, leading to more accurate and efficient decomposition. Our experiments validate that ASVD can compress LLMs by 10%-20% without compromising their reasoning capabilities. Furthermore, ASVD’s compatibility with other LLM compression methods underscores its potential as a versatile tool for LLM deployment.
@@ Appendix A Appendix
Algorithm 1 Binary Search for Truncation Ranks (performance target)
  Input: List of tuples (layer, truncation rank, sensitivity), perplexity target and calibration dataset
  Output: Optimal rank configuration for each layer
  Sort the list by sensitivity in ascending order
  Initialize pointers: $p_{L}=0$, $p_{H}=\text{length of list}-1$
  $p_{M}=\left\lfloor\frac{p_{L}+p_{H}}{2}\right\rfloor$
  while $p_{L}\neq p_{H}$ do
     for each layer in the list do
        Initialize $r=\infty$
        for each tuple in the list starting from $p_{M}$ to the end do
           if tuple’s layer is the same as the current layer then
[…]
           end if
        end for
        if $r=\infty$ then
           Not decompose the layer
        else
           Decompose the layer with truncation rank $r$
        end if
     end for
     Evaluate network performance (perplexity) on calibration dataset
     if perplexity $\leq$ perplexity target then
        $p_{H}=p_{M}$
     else
        $p_{L}=p_{M}+1$
     end if
     Update $p_{M}=\left\lfloor\frac{p_{L}+p_{H}}{2}\right\rfloor$
  end while
A.1 Inference Cost with Decomposed LLMs
@@
A.2 Binary Search for Truncation Ranks
@@
The algorithm of performance target is shown in Algorithm 1. Initially, the low pointer ($p_{L}$) is positioned at the start of the list, while the high pointer ($p_{H}$) is set at the list’s end. The middle pointer ($p_{M}$), as the name suggests, is placed midway between $p_{L}$ and $p_{H}$, calculated as $p_{M}=\left\lfloor\frac{p_{L}+p_{H}}{2}\right\rfloor$. During each iteration of the binary search, we adjust the truncation rank for each layer. Specifically, for a given layer, its truncation rank is set to the smallest rank found to the right of the middle pointer ($p_{M}$) in our list.
@@
Algorithm 2 Binary Search for Truncation Ranks (parameters target)
@@
  Output: Optimal truncation rank configuration for each layer
@@
  $p_{M}=\left\lfloor\frac{p_{L}+p_{H}}{2}\right\rfloor$

### v3

@@ 4.5 Decomposed Network Analysis
We conduct a detailed analysis of the decomposed network. Figure 6 presents the per-type parameters ratio and per-block parameters ratio. Observing the plot, we note that parameters in the MLP components (gate projection, up projection, and down projection) exhibit minimal compression. In MHA, the V projection layer experiences relatively small compression, whereas q projection and k projection can be significantly compressed, indicating redundancy in these components. Turning our attention to the per-block compression ratio, we find that the first layer can undergo substantial compression. In contrast, the compression ratios for the other layers, except for two middle layers, show similar compression rates.
@@ 5 Conclusion
This study presents a training-free approach to compressing Large Language Models (LLMs). We propose Activation-aware Singular Value Decomposition (ASVD) and Sensitivity-based Truncation Rank Searching (STRS), effectively address the challenges posed by activation outliers and varying layer sensitivities. These techniques enable more accurate and efficient decomposition, reducing memory usage and computational demands while maintaining model performance. The successful integration of ASVD into KV cache compression further underscores its potential for broad applicability and substantial impact in real-world scenarios.
@@ Appendix A Appendix
A.1 Impact Statements and Limitations
In this study, we propose a technique that improves the efficiency of Large Language Models (LLMs), making them more accessible. This approach helps to democratize LLMs by lowering deployment costs and hardware barriers, facilitating their use in edge computing. However, it does not mitigate the potential misuse of LLMs by malicious actors.
A.2 Release Safeguards
While ASVD itself does not release new pretrained models, the compression capabilities it provides could enable easier sharing and deployment of powerful models that have risks of misuse. To mitigate risks of misuse, we have implemented access control. Users must agree to terms prohibiting unethical applications.
A.3 Inference Cost with Decomposed LLMs
@@
A.4 Binary Search for Truncation Ranks
@@
The algorithm of performance target: Initially, the low pointer ($p_{L}$) is positioned at the start of the list, while the high pointer ($p_{H}$) is set at the list’s end. The middle pointer ($p_{M}$), as the name suggests, is placed midway between $p_{L}$ and $p_{H}$, calculated as $p_{M}=\left\lfloor\frac{p_{L}+p_{H}}{2}\right\rfloor$. During each iteration of the binary search, we adjust the truncation rank for each layer. Specifically, for a given layer, its truncation rank is set to the smallest rank found to the right of the middle pointer ($p_{M}$) in our list.
@@
1  Output: Optimal truncation rank configuration for each layer
@@
2  $p_{M}=\left\lfloor\frac{p_{L}+p_{H}}{2}\right\rfloor$
@@
Algorithm 1 Binary Search for Truncation Ranks (parameters target)
@@
A.5 Difference with TensorGPT Xu et al. [2023].
In the content of LLM compression via decomposition, the most related work is the concurrent TensorGPT Xu et al. [2023], Zhu et al. [2023], in which the embedding layer of LLMs is compressed through Tensor-Train Decomposition (TTD) Oseledets [2011] in order to store large embeddings in a low-rank tensor format, with much fewer parameters. However, there are several differences between those two methods: (1) Unlike TensorGPT which focuses solely on the token embedding matrix, ASVDaims to compress the entire weight spectrum of LLMs. This holistic approach addresses a more critical aspect of LLM compression, as highlighted in recent studies Lin et al. [2023], Kim et al. [2023]; (2) From the perspective of low-rank decomposition categorization, our method can realize the low-rank decomposition in a rank-adaptive manner, contrasting with the fixed or predetermined ranks used in TensorGPT.
A.6 Empirical Comparison with FWSVD
We also compare ASVD with FWSVD Hsu et al. [2022], which uses Fisher information to weigh the importance of parameters affecting the model prediction. Note that FWSVD is training-required. As shown in Table 4, our method can outperform FWSVD comprehensively.
param ratio 0.95 0.9 0.85 0.8
LLaMA-7b
FWSVD+STRS wiki 5.86 6.32 7.48 10.70
ASVD 5.78 6.09 6.80 8.89
FWSVD+STRS ptb 34.33 38.05 58.75 125.80
ASVD 32.64 37.80 52.11 88.09
LLaMA-2-7b
FWSVD+STRS wiki 5.59 6.12 8.01 13.07
ASVD 5.64 5.93 6.74 8.91
FWSVD+STRS ptb 25.06 36.58 105.53 222.03
ASVD 23.98 32.63 59.84 114.70
A.7 Hyper-parameters Exploration
$\alpha$ 0.1 0.25 0.5 1 2
SVD+STRS 103.39
ASVD abs mean 47.54 37.12 36.89 41.53 43.81
ASVD abs max 52.63 47.17 40.14 41.94 52.55
In our study, we initiate an exploration of hyper-parameters in ASVD, focusing on the activation channel significance metric and the control factor $\alpha$. This exploration is conducted on OPT-125m, a relatively small network that facilitates rapid evaluation.
A.8 Absorbing Singular Values
param ratio weight quant absorbed by UV absorbed by U absorbed by V
0.9 INT6 37.58 39.67 40.62
0.85 INT6 60.44 64.19 61.02